Given a normally distributed population with a mean of 30 and a standard deviation of 8 what percentage of the population will be found above 26?
Answer by stanbon(75887)

Given a normally distributed population with a mean of 30 and a standard deviation of 8 what percentage of the population will be found above 26?
---
Find the z-value of 26.
z(26) = (26-30)/8 = -0.5
---
Then P(x > 26) = P(z > -0.5) = normalcdf(-0.5,100) = 0.6915
---
Cheers,
Stan H.